# “The Diagram is like Guardrails”: Structuring GenAI-assisted Hypotheses Exploration with an Interactive Shared Representation

### v1

@@ 1. Introduction
As Generative AI (GenAI) models continue to advance, there are an increasing number of GenAI-powered tools supporting the execution of specific data analysis tasks, such as creating visualizations. For instance, tools like LIDA (Dibia, 2023) and ChartGPT (Tian et al., 2024) have streamlined the process of generating data visualizations. Researchers also developed interactive task decomposition tool to steer and verify executions in LLM-Assisted data analysis (Kazemitabaar et al., 2024). Meanwhile, there is a growing interest in providing intelligent support for the higher-level conceptual tasks of data analysis. For example, rTisane (Jun, 2024) was designed to support domain assumption checking in data analysis with GenAI. Even at the model level, Large Concept Models (team et al., 2024) have emerged, offering new approaches to higher-level conceptual understanding and generation by operating at a sentence-level semantic representation space.
A direct approach to supporting these higher-level tasks might be to deploy GenAI tools in an autonomous, end-to-end manner, to both produce a large number of possible hypotheses and select appropriate ones to pursue (Zhou et al., 2024; Tong et al., 2024; Si et al., 2024). However, GenAI-driven selection of hypotheses remains an open problem; for instance, recent work has documented challenges for GenAI-powered novelty assessment of hypotheses or scientific papers (Du et al., 2024; Si et al., 2024). A core barrier here is that hypothesis exploration and selection requires judgment with expert domain knowledge of what is more or less interesting or relevant. For instance, in a study on income inequality, a GenAI tool might generate hypotheses about correlations between various demographic factors and income. However, it may not have the nuanced understanding to prioritize hypotheses that are most relevant to current socioeconomic debates or policy considerations. Although such knowledge is pervasive in human society, it often exists in tacit forms, undocumented and unpublished (Ackerman et al., 2013; Cetina, 1999; Rittberg et al., 2018), making it difficult to transfer this decision-making ability to artificial systems.
@@ 2.1. GenAI-Assisted Data Analysis Interfaces: From Task Execution to Conceptual Exploration
Data analysis, as a discipline with mature methodologies, has a vast amount of relevant information available online that has been absorbed by Large Language Models (LLMs), enabling automated execution of specific data analysis subtasks, such as generating data visualizations (Dibia, 2023). LIDA, functioning as a backend engine, can automatically analyze datasets, generate data analysis goals, and produce visualization code and visualizations (Dibia, 2023). Leveraging the data analysis capabilities of LLMs, multiple GenAI-assisted data analysis interfaces have been developed (Ding and Chan, 2024; Vaithilingam et al., 2024; Weidele et al., 2024). While these automated approaches demonstrate the potential of GenAI in data analysis, there is also a need for mixed-initiative (Horvitz, 1999) interaction design to allow users to actively steer and iterate on results produced by GenAI, as opposed to passively accepting the results of an end-to-end process. DynaVis (Vaithilingam et al., 2024) generates data visualizations through a single command and provides an adaptive interface for refining the visualization. Some approaches combine chatbots with graphical user interfaces (GUIs), synchronizing chatbot interactions with corresponding GUI supports for linear, one-way investigation of data (Weidele et al., 2024). In addition to GenAI-assisted systems, Wizard-of-Oz studies have been conducted to observe analysts’ preferences and reactions to GenAI planning assistance (Gu et al., 2024). However, these GenAI interfaces primarily focus on executing analysis procedures with technical precision - “doing things right” - rather than making context-appropriate decision as “doing the right thing”. To create an effective human-AI collaborative interface, in our work, we focus on conceptual subtasks like hypothesis formalization in data analysis (Jun et al., 2022; Jun, 2024). This approach allows human experts to concentrate on decision-making while generative AI implements those decisions, returns results, and suggests next steps. Jun et al. (Jun et al., 2022) proposed the existence of two hypothesis formulation loops in data analysis: the hypothesis refinement loop and the model implementation loop, where the hypothesis refinement loop operates at a higher, conceptual level, indicating the direction of data analysis. Therefore, we build upon their work and select hypothesis exploration as the point of convergence between humans and GenAI, and construct an interface for this intersection.
@@ 2.2. Shared Representation for Data Analysis and Coding
A shared representation between humans and AI constitutes an interface that both can manipulate concurrently in a task-appropriate manner. As AI capabilities progressively advance, systems become increasingly capable of accurately discerning and even anticipating user intentions within the shared representation (Heer, 2019). An example is data wrangling: when a user selects a word in the initial row, the system responds by proposing automatic transformations, such as extracting or truncating text from table cells (Kandel et al., 2011; Guo et al., 2011). In those work, the table serves as a shared representation, allowing the AI to suggest actions and humans to evaluate and refine them. This approach enhances interaction efficiency and accuracy. Alternative methods like textual exchanges or command sequences would make evaluation and iteration more difficult, often obscuring task structure or action outcomes. In AI-assisted code generation, the AI first sketches a code outline based on the user’s prompt before generating the complete code (Chen et al., 2024). Similarly, in frontend code generation, users can sketch the website layout, prompting the AI to generate a corresponding product requirement document, which subsequently informs the code generation process (Zhang et al., 2024). For data analysis, shared representations have evolved from traditional notebooks and dashboards to more sophisticated forms such as canvases (Ding, 2024b), as well as language and block-based interfaces (Cao et al., 2023). An important consideration for shared representations in data analysis is the provision of an overview that contextualizes each instance, facilitating sense-making (Chan et al., 2024). Node-link diagrams, which have been increasingly adopted as a shared representation for Generative AI (Jiang et al., 2023; Huang et al., 2023), show promise as an overview mechanism, elucidating relationships between instances (Chan et al., 2024). In the context of data analysis, node-link diagrams as shared representations align with the multi-thread representation paradigm of multiverse analysis (Dragicevic et al., 2019). Currently, node-link diagrams serve as shared representations for data analytics workflows in systems such as WaitGPT (Xie et al., 2024) and commercialized products such as Altair RapidMiner [^1] https://altair.com/altair-rapidminer and KNIME [^2] https://www.knime.com/. Those applications enable visual coordination of specific subtasks of executing a data analysis workflow, such as moving through data cleaning and executing specific sequences of models. In contrast to the above works, but in line with Jun et al. (Jun, 2024), we are interested in supporting reasoning and exploration of hypotheses, rather than executing data analysis tasks that test these hypotheses.
@@ 3.1. Formative Study
However, this approach introduced new challenges, particularly in terms of information density and display. While F3 hovered over the node, they observed that there were too much information to fit on the screen as the explorating getting deeper, highlighting the limitations of traditional detail-on-demand interaction in this context. To address this issue, participants proposed alternative solutions such as implementing a table and summary view, or potentially opening a new page for more detailed information.
@@ 6.1. Summary of Results
In this paper, we tested the design of a shared representation for GenAI-assisted hypothesis exploration through a design probe (n=22). The majority of our study participants were enthusiastic and excited about this tool and would use it in the future for hypotheses exploration. The proposed ordered node-link diagram serves as a potential shared representation, providing structure while allowing flexibility in exploration. Participants likened it to “guardrails” that focused their thinking without constraining creativity. The diagram facilitated overviews of the hypothesis space and supported frequent backtracking, enabling broader exploration. On average, users created 21.82 hypotheses (SD = 8.78) and reached a maximum depth of 3.86 levels (SD = 1.32), demonstrating exploration patterns including high-level backtracking and branch regeneration. The multimodal information hints, particularly visual hypotheses, helped users judge whether a hypothesis was worth pursuing based on how it might be grounded in the data. The integration of hypotheses and charts reduced context switching effort, while supporting text provided valuable context. These features collectively supported breadth-oriented exploration and depth-oriented iteration in hypothesis formulation.
@@ 6.2.1. Potential System Improvements
The ordered node-link diagram, while generally well-received, had room for improvement according to participant feedback in the interview. Suggestions included enhancing user control over hypothesis exploration, allowing direct manipulation of dataset columns, connecting relevant hypotheses, and showing branched nodes with different user inputs. These improvements could potentially make the system more flexible and informative for users exploring complex data relationships.
@@ 6.2.2. Study Limitation and Future Work
The study has two main limitations that present opportunities for future research. Firstly, the current design lacks between or within-subject A/B testing, which could provide more comparative insights into the effectiveness of different shared representations. To address this, future work should incorporate between-subjects studies to establish statistically significant results. Additionally, the current task setting is confined to a laboratory environment, using a sample 2020 US census dataset and a predetermined analysis intent (“investigate the data from the perspective of income inequality”) that may not fully reflect participants’ real-world data analysis tasks. To enhance ecological validity and broaden the applicability of findings, future research should allow participants to bring their own datasets and define their own analysis tasks or intents. This approach could lead to an “in the wild” deployment, where the system is tested in authentic work environments over extended periods. Such real-world setting would provide deeper insights into how the hypothesis exploration tool integrates into actual data analysis workflows and its long-term impact on conceptual reasoning processes.

### v2

@@ 1. Introduction
As Generative AI (GenAI) models continue to advance, there are an increasing number of GenAI-powered tools supporting the execution of specific data analysis tasks, such as creating visualizations. For instance, tools like LIDA (Dibia, 2023) and ChartGPT (Tian et al., 2024) streamlined the process of generating data visualizations. Researchers also developed interactive task decomposition tools to steer and verify executions in LLM-Assisted data analysis (Kazemitabaar et al., 2024). Meanwhile, there is a growing interest in providing intelligent support for the higher-level conceptual tasks of data analysis. For example, rTisane (Jun, 2024) was designed to support domain assumption checking in data analysis with GenAI. Even at the model level, Large Concept Models (team et al., 2024) have emerged, offering new approaches to higher-level conceptual understanding and generation by operating at a sentence-level semantic representation space.
A direct approach to supporting these higher-level tasks might be to deploy GenAI tools in an autonomous, end-to-end manner, to both produce a large number of possible hypotheses and select appropriate ones to pursue (Zhou et al., 2024; Tong et al., 2024; Si et al., 2024). However, GenAI-driven evaluation and selection of hypotheses remains an open problem such as novelty assessment of hypotheses (Si et al., 2024). A core barrier here is that hypothesis evaluation and selection require judgment with expert domain knowledge of what is more or less interesting or relevant. For instance, in a study on income inequality, a GenAI tool might generate hypotheses about correlations between various demographic factors and income. However, it may not have the nuanced understanding to prioritize hypotheses that are most relevant to current socioeconomic debates or policy considerations. Although such knowledge is pervasive in human society, it often exists in tacit forms, undocumented and unpublished (Ackerman et al., 2013; Cetina, 1999; Rittberg et al., 2018), making it difficult to transfer this decision-making ability to artificial systems.
@@ 2.1. GenAI-Assisted Data Analysis Interfaces: From Task Execution to Conceptual Exploration
Data analysis, as a discipline with mature methodologies, has a vast amount of relevant information available online that has been absorbed by Large Language Models (LLMs), enabling automated execution of specific data analysis subtasks, such as generating data visualizations (Dibia, 2023). LIDA, functioning as a backend engine, can automatically analyze datasets, generate data analysis goals, and produce visualization code and visualizations (Dibia, 2023). Leveraging the data analysis capabilities of LLMs, multiple GenAI-assisted data analysis interfaces have been developed (Ding and Chan, 2024; Vaithilingam et al., 2024; Weidele et al., 2024). While these automated approaches demonstrate the potential of GenAI in data analysis, there is also a need for mixed-initiative (Horvitz, 1999) interaction design to allow users to actively steer and iterate on results produced by GenAI, as opposed to passively accepting the results of an end-to-end process. DynaVis (Vaithilingam et al., 2024) generates data visualizations through a single command and provides an adaptive interface for refining the visualization. Some approaches combine chatbots with graphical user interfaces (GUIs), synchronizing chatbot interactions with corresponding GUI supports for linear, one-way investigation of data (Weidele et al., 2024). However, these GenAI interfaces primarily focus on executing analysis procedures with technical precision - “doing things right” - rather than making context-appropriate decision as “doing the right thing”. Some prior work suggests desire paths for GenAI-assisted approaches that extend beyond execution, to assist with more conceptual tasks like planning data analysis (Gu et al., 2024). Other conceptual tasks may include exploratory data analysis (EDA) to probe and understand structures and patterns in the data to inform planning and hypothesizing. In this work, we build specifically on Jun et al’s modeling of hypothesis formalization as a key subtask of data analysis: iterating over the space of conceptual hypotheses in conversation with properties of the data to formalize hypotheses into statistical models and procedures that can then be executed (Jun et al., 2022; Jun, 2024). This subtask of hypothesis formalization differs from EDA in that it explicitly focuses on exploring and formalizing conceptual hypotheses (which can often be stated in natural language) rather than broadly probing for structures and patterns with visualizations of data distributions and properties.
@@ 2.2. Shared Representation for Data Analysis and Coding
A shared representation between humans and AI constitutes an interface that both can manipulate concurrently in a task-appropriate manner. As AI capabilities progressively advance, systems become increasingly capable of accurately discerning and even anticipating user intentions within the shared representation (Heer, 2019). An example is data wrangling: when a user selects a word in the initial row, the system responds by proposing automatic transformations, such as extracting or truncating text from table cells (Kandel et al., 2011; Guo et al., 2011). In those work, the table serves as a shared representation, allowing the AI to suggest actions and humans to evaluate and refine them. This approach enhances interaction efficiency and accuracy. Alternative methods like textual exchanges or command sequences would make evaluation and iteration more difficult, often obscuring task structure or action outcomes. In AI-assisted code generation, the AI first sketches a code outline based on the user’s prompt before generating the complete code (Chen et al., 2024). Similarly, in frontend code generation, users can sketch the website layout, prompting the AI to generate a corresponding product requirement document, which subsequently informs the code generation process (Zhang et al., 2024). For data analysis, shared representations have evolved from traditional notebooks and dashboards to more sophisticated forms such as canvases (Ding, 2024b), as well as language and block-based interfaces (Cao et al., 2023). An important consideration for shared representations in data analysis is the provision of an overview that contextualizes each instance, facilitating sense-making (Chan et al., 2024). Node-link diagrams, which have been increasingly adopted as a shared representation for Generative AI (Jiang et al., 2023; Huang et al., 2023), show promise as an overview mechanism, elucidating relationships between instances (Chan et al., 2024). Currently, node-link diagrams serve as shared representations for data analytics workflows in systems such as WaitGPT (Xie et al., 2024) and commercialized products such as Altair RapidMiner [^1] https://altair.com/altair-rapidminer and KNIME [^2] https://www.knime.com/. Those applications enable visual coordination of specific subtasks of executing a data analysis workflow, such as moving through data cleaning and executing specific sequences of models. We build on these prior explorations to investigate the potential of a node-link diagram pattern as a shared representation for supporting GenAI-assisted hypothesis exploration. We hypothesize that the structure of node-link diagrams could facilitate their use as shared representations given their alignment with the multi-thread representation paradigm of multiverse analysis (Dragicevic et al., 2019).
@@ 3.1. Formative Study
However, this approach introduced new challenges, particularly in terms of information density and display. While F3 hovered over the node at the bottom, they observed that there was too much information to fit on the screen as the exploration got deeper, highlighting the limitations of traditional detail-on-demand interaction in this context. To address this issue, participants proposed alternative solutions such as implementing a table and summary view, or potentially opening a new page for more detailed information.
@@ 6.1. Summary of Results
In this paper, we tested the design of a shared representation for GenAI-assisted hypothesis exploration with a design probe (n=22). The majority of our study participants were enthusiastic and excited about this tool and would use it in the future for hypotheses exploration. The proposed ordered node-link diagram serves as a potential shared representation, providing structure while allowing flexibility in exploration. Participants likened it to “guardrails” that focused their thinking without constraining creativity. The diagram facilitated overviews of the hypothesis space and supported frequent backtracking, enabling broader exploration. On average, users created 21.82 hypotheses (SD = 8.78) and reached a maximum depth of 3.86 levels (SD = 1.32), demonstrating exploration patterns including high-level backtracking and branch regeneration. The multimodal information hints, particularly visual hypotheses, helped users judge whether a hypothesis was worth pursuing based on how it might be grounded in the data. The integration of hypotheses and charts reduced context switching effort, while supporting text provided valuable context. These features collectively supported breadth-oriented exploration and depth-oriented iteration in hypothesis formulation.
@@ 6.2.1. Potential System Improvements
The ordered node-link diagram, while generally well-received, had room for improvement according to participant feedback in the interview. Suggestions included enhancing user control over hypothesis exploration, allowing direct manipulation of dataset columns, connecting relevant hypotheses, and showing branched nodes with different user inputs.
Another limitation arises from the scalability of node-link diagrams: as more hypotheses are explored, the diagram’s complexity increases rapidly. In the short time period of our study, participants explored on the order of 20-30 hypothesis nodes, on average: it is possible that, accounting for variations in screen real estate across devices, that this is close to the practical upper bound for what can be visually managed. It is also possible, however, that in the context of larger data analysis workflows, the scale of exploration might not extend much more beyond tens of hypothesis nodes. Future work should systematically evaluate hypothesis exploration using node-link diagrams in more ecologically valid settings, such as with problems of varying complexity, or over longer time periods, to better understand these scalability limitations, and explore how augmenting the node link diagram with previously studied techniques, such as dynamic rescaling and filtering/aggregation of branches (Plaisant et al., 2002) or partial overviews (Lee et al., 2006), might overcome these limitations.
@@
Finally, we envision our system functioning as a component in data analysis workflows, to be used during conceptual hypothesis exploration, rather than replacing entire existing workflows, such as EDA. How our prototype system might interact with other components of data analysis workflows remains an open question. Perhaps more specific EDA visualizations might enhance the information hints in our interface, and selected hypotheses from our system could be ”exported” to statistical model planning systems like rTisane for further refinement (Jun, 2024).
@@ 6.2.2. Study Limitation and Future Work
The study has three main limitations that present opportunities for future research. First, our research questions and study design here focused on qualitative understanding of the user experience; thus, our data cannot speak to quantitative or comparative effects of our system on the quality or creativity of hypothesis exploration. To address this, future work should incorporate direct comparisons to baseline or other systems in appropriate between or within-subjects studies; these studies could compare process-based metrics (e.g., number/diversity/depth of hypotheses explored) as well outcome-based metrics (probability of insights, overall quality/novelty of hypotheses).
Additionally, the current task setting is confined to a laboratory environment, using a sample 2020 US census dataset and a predetermined analysis intent (“investigate the data from the perspective of income inequality”) that may not fully reflect participants’ real-world data analysis tasks. To enhance ecological validity and broaden the applicability of findings, future research should allow participants to bring their own datasets and define their own analysis tasks or intents. This approach could lead to an “in the wild” deployment, where the system is tested in authentic work environments over extended periods. Such real-world setting would provide deeper insights into how the hypothesis exploration tool integrates into actual data analysis workflows and its long-term impact on conceptual reasoning processes.
Third, our system builds on prior work that benchmarks the quality, novelty, and other dimensions of AI-generated hypotheses from LLMs comparable to those we used (Si et al., 2024; Baek et al., 2025). Therefore, we assumed a baseline level of quality and did not independently conduct a comprehensive evaluation of the generated hypotheses. While Table 2 presents hypothesis examples, we were not yet in a position to make claims about the overall quality of the generated content. Future work could evaluate these hypotheses across multiple dimensions, including feasibility, potential impact, and novelty.